Present Value for Various Compounding Periods

Find the present value of $500 due in the future under each of the following conditions. Do not round intermediate calculations. Round your answers to the nearest cent.

  1. 4% nominal rate, semiannual compounding, discounted back 5 years
    $  
  2. 4% nominal rate, quarterly compounding, discounted back 5 years
    $  
  3. 4% nominal rate, monthly compounding, discounted back 1 year
    $  
Answer #1

a)

Future Value (FV)

$500

Interest Rate - Semi-annual (i)

4percent / 2

= 2 percent (half-yearly)

Number of Periods (n)

5 years x 2

= 10 periods

According to Problem:

PV = FV/(1+i)

PV = 500/(1+2/10010

PV = 410.17

Therefore, PRESENT VALUE = $410.17

b)

Future Value (FV)

$500

Interest Rate – Quarterly Compounded (i)

4percent / 4

= 1 percent (per quarter)

Number of Periods (n)

5 years x 4

= 20 periods

According to Problem:

PV = FV/(1+i)

PV = 500/(1+1/10020

PV = 409.77

Therefore, PRESENT VALUE = $409.77

c)

Future Value (FV)

$500

Interest Rate – monthly compounded (i)

4percent / 12

= 0.33% percent (per month)

Number of Periods (n)

1 year x 12

= 12 periods

According to Problem:

PV = FV/(1+i)

PV = 500/(1+0.33/100)12

PV = 480.62

Therefore, PRESENT VALUE = $480.62